OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope?Point: (–8, 3); Slope: 6 A. y + 3 = 6x – 48 B. y + 3 = 6(x + 8) C. y – 3 = 6(x – 8) D. y – 3 = 6(x + 8) B?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Yes, you're correct!

OpenStudy (anonymous):

so y + 3 = 6(x + 8) ? just making sure lol

OpenStudy (anonymous):

Oh, wait. You have your numbers correctly but not your signs..

OpenStudy (anonymous):

y – 3 = 6(x – 8) ?

OpenStudy (anonymous):

Incorrect. A negative + negative = negative

OpenStudy (anonymous):

D! y – 3 = 6(x + 8) right?